OpenStudy (anonymous):

(person who helps gets medal) Indicate in standard form the equation of the line passing through the given point and having the given slope. R(4, 0), m = 5

OpenStudy (anonymous):

@Vocaloid can you help

Vocaloid (vocaloid):

first use the point-slope equation

Vocaloid (vocaloid):

y-y1 = m(x-x1) where m = slope and (x1,y1) = the given point

OpenStudy (anonymous):